# CMMU: A Benchmark for Chinese Multi-modal Multi-type Question Understanding and Reasoning

### v1

@@ Abstract
Multi-modal large language models(MLLMs) have achieved remarkable progress and demonstrated powerful knowledge comprehension and reasoning abilities. However, the mastery of domain-specific knowledge, which is essential for evaluating the intelligence of MLLMs, continues to be a challenge. Current multi-modal benchmarks for domain-specific knowledge concentrate on multiple-choice questions and are predominantly available in English, which imposes limitations on the comprehensiveness of the evaluation. To this end, we introduce CMMU, a novel benchmark for multi-modal and multi-type question understanding and reasoning in Chinese. CMMU consists of 3,603 questions in 7 subjects, covering knowledge from primary to high school. The questions can be categorized into 3 types: multiple-choice, multiple-response, and fill-in-the-blank, bringing greater challenges to MLLMs. In addition, we propose a rigorous evaluation strategy called ShiftCheck for assessing multiple-choice questions. The strategy aims to reduce position bias, minimize the influence of randomness on correctness, and perform a quantitative analysis of position bias. We evaluate seven open-source MLLMs along with GPT4-V, Gemini-Pro, and Qwen-VL-Plus. The results demonstrate that CMMU poses a significant challenge to the recent MLLMs.
Figure 1: Some examples in MMCU. We provide Chinese examples and their corresponding English translations.
@@ 1 Introduction
To bridge the dataset gap, we introduce a novel benchmark, CMMU, for multi-modal and muli-type question understanding and reasoning in Chinese. CMMU encompasses multi-modal content across 7 subjects. Every question requires the model to combine image and text content to generate a comprehensive response. While CMMU shares similarities with datasets like ScienceQA and M3Exam Zhang et al. (2023), it offers a broader range of question types. Previous datasets only have multiple-choice questions, while CMMU offers a wider variety of question types, including multiple-choice, multiple-response, and fill-in-the-blank questions, which poses a more significant challenge to the comprehension abilities of MLLM. In addition, to mitigate the position biasZheng et al. (2023) in LLM and ensure genuine correctness rather than guessing, inspired by CircularEval Liu et al. (2023c), we adopt a ShiftCheck approach for multiple-choice questions. Specifically, we cycle through the position of options to ensure that the answer can appear at any position with equal probability, aim at reducing position bias, minimizing the influence of randomness on correctness, and performing a quantitative analysis of position bias. We evaluate 11 models using the CMMU benchmark, and the results indicate that CMMU presents a significant challenge to current MLLMs.
Figure 2: Distribution of the number of questions in each subject and grades.
@@
  • We propose the ShiftCheck strategy, which is designed to rigorously evaluate multiple-choice questions and conduct a quantitative analysis of position bias in MLLMs.
@@ 2.1 Multi-modal Benchmarks
There are already many datasets for evaluating the multi-modal capabilities of models. For instance, datasets like VQAv2 Goyal et al. (2017), TDIUC Kafle and Kanan (2017), TextVQA Singh et al. (2019), and GQA Hudson and Manning (2019) are used in visual question answering tasks, while COCO Lin et al. (2014), NoCaps Agrawal et al. (2019), and Flickr30K Plummer et al. (2015) are employed in image captioning tasks. Additionally, Visual7w Zhu et al. (2016) and RefCOCO Kazemzadeh et al. (2014) are commonly utilized for visual grounding purposes. With the rapid development of multi-modal large language models, researchers have achieved good results on these datasets. We require more extensive data to evaluate MLLMs, and there have been recent studies evaluating models from various perspectives. LVLM-eHub Xu et al. (2023) collects 47 existing benchmarks and evaluates 6 types of capabilities of MLLMs, however, it does not create any new benchmarks. MME Fu et al. (2023) comprehensively measures the perception and cognition abilities of models. However, its question types are simplistic, merely requiring yes or no responses. MMBench Liu et al. (2023c) and SEED-Bench Li et al. (2023b, a) contain many multiple-choice questions covering various ability dimensions, but these datasets mainly consist of common-sense questions and do not require lots of domain-specific knowledge and complex reasoning. To enhance the evaluation of domain-specific knowledge, ScienceQA Lu et al. (2022) was introduced. This dataset encompasses a wide range of science topics from elementary and high school curricula. MMMU Yue et al. (2023) is designed to evaluate college-level subject knowledge, questions of CMMU are collected from college exams and textbooks, and many of them require expert-level skills. M3Exam Zhang et al. (2023) is a multilingual and multi-modal benchmark designed to evaluate domain knowledge and problem-solving skills, and it spans seven languages. However, less than one-third of the questions include images.
@@ 2.2 Multi-modal Large Language Models
Average explanation length 311.67
@@ 3 The CMMU Benchmark
Figure 5: A demonstration of the Circular Evaluation in ShiftCheck.
@@ 3.1 Data Pre-process
Data Cleaning. We manually review the questions, filtering out images that are blurry, low-quality, or have a resolution less than 50$\times$50 dpi, eliminating questions that are incorrectly parsed, and correcting mistakes made during the automatic conversion to LaTeX. Furthermore, experienced teachers consider the depth of knowledge and the complexity of question-solving methods to categorize each question into two levels: normal and hard.
@@ 3.2 Data Distribution
CMMU has a total of 3603 questions, divided into validation set and test set, with 1,800 and 1,803 questions respectively. The validation set will be open source to the community. As shown in Figure 2, the benchmark contains multi-modal content from middle and high school across 7 types of subjects, while primary school only contains math. The ratio of normal and hard questions are 8:2. Over 70% of the questions have detailed answer explanations, with an average length of 311 characters for each analysis. The detailed statistics are shown in Table 1. We translate the original questions into English and analyze the distribution of their first three words. As shown in Figure 3, the questions have diverse formats and are relevant to images or diagrams.
@@ 4.1 ShiftCheck Strategy
To address the above problems, we propose the ShiftCheck Strategy including two stages. Firstly, following the Circular Evaluation Liu et al. (2023c), we cyclically change the positions of the options and let the model answer questions. Subsequently, we calculate metrics to quantify position bias. We will describe the whole process in detail.
Circular Evaluation For a multiple-choice question with $k$ options, we perform a right circular shift on the options. For example, if the original order of the options is $ABCD$, then after one shift, the order will change to $DABC$. A detailed example is provided in Figure 5. Given a question $Q$ with $k$ possible options, we generate $k$ distinct shifted-option questions, denoted as $Q_{i}$, $i\in[0,k]$. Each $Q_{i}$ is then evaluated by the MLLMs to generate the corresponding answers $A_{i}$. We consider the model to have sufficient knowledge to answer the question $Q$ only if all of $A_{i}$ are correct, in which case the accuracy score of $Q$ is 1, otherwise it is 0.
@@
Intern-XComposer-7b 14.0 16.6 6.5 31.82 22.82 19.39 34.27 6.83 23.06 16.07
Qwen-VL-Plus 17.26 20.29 7.0 52.64 37.46 27.18 48.82 12.67 33.8 21.36
@@
Position Bias Measurement Conceptually, an unbiased model assigns equal probability to each option. Under the shifted-option setting, if the probability of each option is not equal, it indicates that the model has a bias towards a certain option. Considering this, we define the $BiasRate$ as follows:
Questions that are completely answered correctly in Circular Evaluation do not reflect position bias, so we just focus on the incorrectly answered questions. If there are $M$ incorrectly answered questions with $n$ options for each, there will be a total of $m*n$ answers combination. We count the occurrences $S_{o}$ of each option $o$ and then calculate the probability $P_{o}=\frac{S_{o}}{m*n},o\in\{A,B,C,...\}$. And then we define the BiasRate as the variance of $P$, the formula is $BiasRate=\sigma^{2}(P)$. The larger the $BiasRate$, the greater the positional bias of the model.
@@ 4.2 Evaluations on Different Question Types
Evaluation on Multiple-choice Question: We adopt the ShiftCheck Strategy as mentioned in section 4.1, which allows us to analyze both the accuracy and the $BiasRate$.
@@ 5.1 Models
We evaluate the performance of various MLLMs on the CMMU benchmark, including both closed-source and open-source models. The closed-source models are evaluated by using their official API, while open-source models are evaluated by running inferences on NVIDIA A100 GPUs. For the closed-source models, we select state-of-the-art models like GPT-4V, Gemini-Pro. We also choose Qwen-VL-Plus, which performs well on Chinese datasets. For the open-source models, model sizes vary from 7b to 13b, including LLava-1.5-13b, CogVLM-7b, Qwen-VL-Chat-7b, InstructBLIP-13b, Intern-XComposer-7b, mPLUG-Owl2-7b and ShareGPT4V-7b.
@@ 5.3.2 Results on Different Question Types
Figure 6 shows the overall results of different models in subjects and grades, and the detailed statistics are shown in Table 4. We can see that there is a significant subject bias in all models. Subjects such as politics and history rely on the knowledge reservoir and have higher accuracy than subjects such as physics, math, and biology, which require computation and reasoning. When comparing different grades, the results show that MLLMs generally perform better on middle school questions than high school ones, suggesting that more complex knowledge presents a more significant challenge for these models.
@@ 5.3.3 Position Bias Analysis
We employ the ShiftCheck Strategy for quantitative analysis of position bias. As shown in Figure 7, most models have a positional preference for one or two specific options. An interesting finding is that, although these positional preferences are inconsistent across models, none of them choose Option $D$ as their most preferred choice. By analyzing the $BiasRate$, we find that superior models, such as GPT-4V, tend to have a relatively lower $BiasRate$.
@@ 5.3.4 Case Study with CoT Prompts
To further analyze the performance of models using Chain of Thought(CoT), we change the prompt of MCQ to “Please analyze the question step by step and eventually provide a single correct option letter. (This is a multiple-choice question.)” Then, we choose GPT-4V, which has a strong ability in instruction-following, to answer 500 randomly selected MCQs. We identify three common mistake types in the model outputs: image misunderstanding, misleading reasoning, and question misunderstanding, with proportions of 27.48%, 35.41%, and 13.03%, respectively. Cases in Figure 8 show the above common mistakes respectively: In Question 1, the model fails to identify the route A in the image correctly. In Question 2, the model thinks that the kinetic energy during free fall is transformed into gravitational potential energy, leading to an incorrect reasoning result. In Question 3, the model misunderstands the question and produces a hallucination that the overlap area is 1. Although it guesses the answer correctly, it cannot pass the ShiftCheck Strategy. All bad cases demonstrate that even one of the most advanced MLLMs cannot perfectly solve questions of CMMU, highlighting both the potential and challenges of the benchmark.
@@ 6 Conclusion and Future Work
In conclusion, our work introduces a novel benchmark named CMMU to evaluate the multi-modal and multi-type question understanding and reasoning abilities of MLLMs in Chinese. Unlike existing benchmarks focusing on multiple-choice questions, CMMU offers a more comprehensive evaluation by incorporating a broader question type, including MCQ, MRQ, and FBQ. We also propose the ShiftCheck approach to quantify the position bias of the model, minimize the impact of randomness and ensure evaluation correctness. The evaluation results contribute to a deeper understanding of current MLLMs in the context of diverse and complex question formats. In future work, we will consider enriching the problem types of CMMU and expanding the number of problems to increase the challenge of the benchmark further.

### v3

@@ Abstract
Multi-modal large language models(MLLMs) have achieved remarkable progress and demonstrated powerful knowledge comprehension and reasoning abilities. However, the mastery of domain-specific knowledge, which is essential for evaluating the intelligence of MLLMs, continues to be a challenge. Current multi-modal benchmarks for domain-specific knowledge concentrate on multiple-choice questions and are predominantly available in English, which imposes limitations on the comprehensiveness of the evaluation. To this end, we introduce CMMU, a novel benchmark for multi-modal and multi-type question understanding and reasoning in Chinese. CMMU consists of 3,603 questions in 7 subjects, covering knowledge from primary to high school. The questions can be categorized into 3 types: multiple-choice, multiple-response, and fill-in-the-blank, bringing greater challenges to MLLMs. In addition, we propose an evaluation strategy called Positional Error Variance for assessing multiple-choice questions. The strategy aims to perform a quantitative analysis of position bias. We evaluate seven open-source MLLMs along with GPT4-V, Gemini-Pro, and Qwen-VL-Plus. The results demonstrate that CMMU poses a significant challenge to the recent MLLMs. The data and code are available at https://github.com/FlagOpen/CMMU.
Figure 1: Some examples in CMMU. We provide Chinese examples and their corresponding English translations.
@@ 1 Introduction
To bridge the dataset gap, we introduce a novel benchmark, CMMU, for multi-modal and muli-type question understanding and reasoning in Chinese. CMMU encompasses multi-modal content across 7 subjects. Every question requires the model to combine image and text content to generate a comprehensive response. While CMMU shares similarities with datasets like ScienceQA and M3Exam Zhang et al. (2023), it offers a broader range of question types. Previous datasets only have multiple-choice questions, while CMMU offers a wider variety of question types, including multiple-choice, multiple-response, and fill-in-the-blank questions, which poses a more significant challenge to the comprehension abilities of MLLM. In addition, to mitigate the position biasZheng et al. (2023) in LLM and ensure genuine correctness rather than guessing, inspired by CircularEval Liu et al. (2023c), we adopt a Positional Error Variance approach to measure the position bias for multiple-choice question. Specifically, we cycle through the position of options to ensure that the answer can appear at any position with equal probability, which is the same as CircularEval, aim at reducing position bias, minimizing the influence of randomness on correctness. Then we produce Positional Error Variance, a quantitative analysis to measure position bias. We evaluate 11 models using the CMMU benchmark, and the results indicate that CMMU presents a significant challenge to current MLLMs.
Figure 2: Distribution of questions in subjects and grades.
@@
  • We propose Positional Error Variance, which is designed to conduct a quantitative analysis of position bias in MLLMs.
@@ 2.1 Multi-modal Benchmarks
With the development of large language models (LLMs), there is a growing emphasis in research communities on assessing the capabilities of LLM such as HELMLiang et al. (2022), CLEVA Li et al. (2023d), especially their multimodal understanding abilities. Datasets like VQAv2 Goyal et al. (2017), TDIUC Kafle and Kanan (2017), TextVQA Singh et al. (2019) and GQA Hudson and Manning (2019) are used in visual question answering tasks, while COCO Lin et al. (2014), NoCaps Agrawal et al. (2019), and Flickr30K Plummer et al. (2015) are employed in image captioning tasks. Additionally, Visual7w Zhu et al. (2016) and RefCOCO Kazemzadeh et al. (2014) are commonly utilized for visual grounding purposes. With the rapid development of multi-modal large language models, researchers have achieved good results on these datasets. We require more extensive data to evaluate MLLMs, and there have been recent studies evaluating models from various perspectives. LVLM-eHub Xu et al. (2023) collects 47 existing benchmarks and evaluates 6 types of capabilities of MLLMs, however, it does not create any new benchmarks. MME Fu et al. (2023) comprehensively measures the perception and cognition abilities of models. However, its question types are simplistic, merely requiring yes or no responses. MMBench Liu et al. (2023c) and SEED-Bench Li et al. (2023b, a) contain many multiple-choice questions covering various ability dimensions, but these datasets mainly consist of common-sense questions and do not require lots of domain-specific knowledge and complex reasoning. To enhance the evaluation of domain-specific knowledge, ScienceQA Lu et al. (2022) was introduced. This dataset encompasses a wide range of science topics from elementary and high school curricula. MMMU Yue et al. (2023) is designed to evaluate college-level subject knowledge, questions of CMMU are collected from college exams and textbooks, and many of them require expert-level skills. BenchLMM Cai et al. (2023) assesses MLMMs in three distinct styles: artistic style, sensor style, and application style. M3Exam Zhang et al. (2023) is a multilingual and multi-modal benchmark designed to evaluate domain knowledge and problem-solving skills, and it spans seven languages. However, less than one-third of the questions include images.
@@ 2.2 Multi-modal Large Language Models
Average answer explanation length 311.67
@@ 3 The CMMU Benchmark
Figure 5: A demonstration of the CircularEval.
@@ 3.1 Data Pre-process
Data Cleaning. We manually review the questions, filtering out images that are blurry, low-quality, or have a resolution less than 50$\times$50 dpi, eliminating questions that are incorrectly parsed, and correcting mistakes made during the automatic conversion to LaTeX. Furthermore, experienced teachers consider the depth of knowledge and the complexity of question-solving methods to categorize each question into two levels: normal and hard based on their experience and subjective judgment.
@@ 3.2 Data Distribution
CMMU has a total of 3603 questions, divided into validation set and test set, with 1,800 and 1,803 questions respectively. The validation set is open source to the community. As shown in Figure 2, the benchmark contains multi-modal content from middle and high school across 7 types of subjects, while primary school only contains math. The ratio of normal and hard questions are 8:2. Over 70% of the questions have detailed answer explanations, with an average length of 311 characters for each analysis. The detailed statistics are shown in Table 1. We translate the original questions into English and analyze the distribution of their first three words. As shown in Figure 3, the questions have diverse formats and are relevant to images or diagrams.
@@ 4 Evaluation
To address the above problems, we propose Positional Error Variance. Firstly, following the CircularEval Liu et al. (2023c), we cyclically change the positions of the options and let the model answer questions. Subsequently, we calculate metrics to quantify position bias. We will describe the whole process in detail.
4.1 CircularEval
For a multiple-choice question with $k$ options, we perform a right circular shift on the options. For example, if the original order of the options is $ABCD$, then after one shift, the order will change to $DABC$. A detailed example is provided in Figure 5. Given a question $Q$ with $k$ possible options, we generate $k$ distinct shifted-option questions, denoted as $Q_{i}$, $i\in[0,k]$. Each $Q_{i}$ is then evaluated by the MLLMs to generate the corresponding answers $A_{i}$. We consider the model to have sufficient knowledge to answer the question $Q$ only if all of $A_{i}$ are correct, in which case the accuracy score of $Q$ is 1, otherwise it is 0.
@@ 4.1 CircularEval
Intern-XComposer-7b 12.92 15.82 6.02 30.14 22.67 18.65 33.8 6.48 22.07 15.59
Qwen-VL-Plus 17.57 22.12 6.88 52.88 37.76 29.15 49.29 11.99 34.79 22.14
@@
4.2 Positional Error Variance
Conceptually, an unbiased model assigns equal probability to each option. Under the shifted-option setting, if the probability of each option is not equal, it indicates that the model has a bias towards a certain option. Considering this, we define the $BiasRate$ as follows:
Questions that are completely answered correctly in CircularEval do not reflect position bias, so we just focus on the incorrectly answered questions. If there are $M$ incorrectly answered questions with $n$ options for each, there will be a total of $m*n$ answers combination. We count the occurrences $S_{o}$ of each option $o$ and then calculate the probability $P_{o}=\frac{S_{o}}{m*n},o\in\{A,B,C,...\}$. And then we define the BiasRate as the variance of $P$, the formula is $BiasRate=\sigma^{2}(P)$. The larger the $BiasRate$, the greater the positional bias of the model.
@@ 4.3 Evaluations on Different Question Types
Evaluation on Multiple-choice Question: We adopt CircularEval and Positional Error Variance in section 4, which allows us to analyze both the accuracy and the $BiasRate$.
@@ 5.1 Models
We evaluate the performance of various MLLMs, including both closed-source and open-source models. The closed-source models are evaluated by using their official API, while open-source models are evaluated by running inferences on NVIDIA A100 GPUs. For the closed-source models, we select state-of-the-art models like GPT-4V, Gemini-Pro. We also choose Qwen-VL-Plus, which performs well on Chinese datasets. For the open-source models, model sizes vary from 7b to 13b, including LLava-1.5-13b, CogVLM-7b, InstructBLIP-13b, Qwen-VL-Chat-7b, Intern-XComposer-7b, mPLUG-Owl2-7b and ShareGPT4V-7b.
@@ 5.3.2 Results on Different Question Types
Figure 6 shows the overall results of different models in subjects and grades, and the detailed statistics are shown in Table 4. We can see that there is a significant subject bias in all models. Subjects such as politics and history rely on the knowledge reservoir have higher accuracy than subjects such as physics, math which require computation and reasoning. When comparing different grades, the results show that MLLMs generally perform better on middle school questions than high school ones, suggesting that more complex knowledge presents a more significant challenge for these models.
@@ 5.3.3 Position Bias Analysis
We employ the Positional Error Variance for quantitative analysis of position bias. As shown in Figure 7, most models have a positional preference for one or two specific options. An interesting finding is that, although these positional preferences are inconsistent across models, none of them choose Option $D$ as their most preferred choice. By analyzing the $BiasRate$, we find that superior models, such as GPT-4V, tend to have a relatively lower $BiasRate$.
@@ 5.3.4 Case Study with CoT Prompts
To further analyze the performance of models using Chain of Thought(CoT), we change the prompt of MCQ to “Please analyze the question step by step and eventually provide a single correct option letter. (This is a multiple-choice question.)” Then, we choose GPT-4V, which has a strong ability in instruction-following, to answer 500 randomly selected MCQs. We identify three common mistake types in the model outputs: image misunderstanding, misleading reasoning, and question misunderstanding, with proportions of 27.48%, 35.41%, and 13.03%, respectively. Cases in Figure 8 show the above common mistakes respectively: In Question 1, the model fails to identify the route A in the image correctly. In Question 2, the model thinks that the kinetic energy during free fall is transformed into gravitational potential energy, leading to an incorrect reasoning result. In Question 3, the model misunderstands the question and produces a hallucination that the overlap area is 1. Although it guesses the answer correctly, it cannot pass the CircularEval. All bad cases demonstrate that even one of the most advanced MLLMs cannot perfectly solve questions of CMMU, highlighting both the potential and challenges of the benchmark.
@@ 6 Conclusion and Future Work
In conclusion, our work introduces a novel benchmark named CMMU to evaluate the multi-modal and multi-type question understanding and reasoning abilities of MLLMs in Chinese. Unlike existing benchmarks focusing on multiple-choice questions, CMMU offers a more comprehensive evaluation by incorporating a broader question type, including MCQ, MRQ, and FBQ. We also propose Positional Error Variance to quantify the position bias of the model. The evaluation results contribute to a deeper understanding of current MLLMs in the context of diverse and complex question formats. In future work, we will consider enriching the problem types of CMMU and expanding the number of problems to increase the challenge of the benchmark further.